daq/hpss status

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Thu Aug 23 2001 - 07:21:34 EDT

  • Next message: Apache: "Shift report 20010823 00:00-08:00"

    I should like to clarify a few points 
    (add this to logBook please)
    a) Though not clear at time the reason for the failing re-boot of opus was not hardware but the interaction with the network
       (gigabit to RCF)
    b) The data runs should have been chekced as write to HPSS, since the will prevent the data files from being erased, and transmitted
        once the links get restored (hopefully today). Once the spool area is full data taking would stop, but could resume by writting to
        DLT-tape. This later procedure is not described, and have not been tested for a long time, but should work.
    
    
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b30 : Thu Aug 23 2001 - 07:13:52 EDT